OpenStudy (jackellyn):

Find a basis for the following subspace of R^4. All vectors of the form (a+c, a-b, b+c, -a+b).

OpenStudy (amistre64):

i would set it up so that they all have an a,b,c component

OpenStudy (jackellyn):

So far I have \[\left(\begin{matrix}a+c \\ a-b \\ b+c \\ -a+b\end{matrix}\right)=a \left(\begin{matrix}1 \\ 1 \\0 \\-1\end{matrix}\right)+b \left(\begin{matrix}0 \\ -1 \\1 \\1\end{matrix}\right)+c \left(\begin{matrix}1 \\ 0 \\1 \\0\end{matrix}\right)\]

OpenStudy (amistre64):

1a + 0b + 1c 1a - 1b + 0c 0a + 1b + 1c -1a + 1b + 0c then strip the a,b,c parts and row reduce 1 0 1 1 -1 0 0 1 1 -1 1 0

OpenStudy (amistre64):

a basis, if i recall, is a most efficient span; so we would have to determine the pivot columns

OpenStudy (amistre64):

1 0 1 1 -1 0 0 1 1 -1 1 0 1 0 1 -1 1 0 0 1 1 0 1 1 1 0 1 0 1 1 0 1 1 0 1 1 1 0 1 0 1 1 0 0 0 0 0 0 looks like we have a free variable; and 2 pivot columns

OpenStudy (amistre64):

basis = [1 1 0 -1], [0,-1,1,1] might be it

OpenStudy (amistre64):

opps, got comma happy :) those are spose to be column vectors basis = [1 1 0 -1], [0 -1 1 1] might be it

OpenStudy (anonymous):

@amistre64 agreed